OpenStudy (anonymous):

Please Help:Find the surface area of the solid. The pyramids are regular and the prisms, cylinders, and cones are right. Round results to one decimal place..Please Help

OpenStudy (anonymous):

i got 30ft for the first one

OpenStudy (anonymous):

This might sound redundant, but just find the area of the shapes on the surface and add them all up. So for the first one you have a pyramid, so find the area of one triangle and multiply it by the number of sides/surfaces of the pyramid to get the surface area of the pyramid. Do the same for the cube and add everything up. Pyramid should be 30ft squared and the cube is 45 ft squared for a total of 75 ft squared of surface area. Apply the same method for the other shapes.

OpenStudy (anonymous):

so the first one is 75 ft squared?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so for the second one i got 107.39 for the cone

OpenStudy (anonymous):

and for the bottom part i got 27.14.. so than i add those two together?

OpenStudy (anonymous):

Surface area of a cone is pi*r*s + the base. Where r is the radius and s is the length of the side of the cone. So we have 3.14*3*8.4=79.168. Then you need to add the bottom cylinder, which is circumference times height giving 2pi*r*h which is 96.133. Then lastly add the surface area of the base of the cylinder which is just pi*r^2 giving 28.274. Add all them up and you have your answer in cm squared.

OpenStudy (anonymous):

the formula i was showed to use is cone: S=pi*r^2+pi*r*l and for the cylinder: S=2*pi*r^2+pi*r*h

OpenStudy (anonymous):

The cone formula is correct. Your pi*r*l is the same as my pi*r*s. The other part is accounting for the base of the cone, which technically isn't a surface. The cylinder formula is 2*pi*r*h for the sides and then you do pi*r^2 for the base.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i got 211.95cm^2

OpenStudy (anonymous):

I think when I did it I got 203.6 cm^2.

OpenStudy (anonymous):

ok